OpenStudy (anonymous):

Square root questions!! medal and fan if you can help:

OpenStudy (anonymous):

ok whats your question

OpenStudy (anonymous):

wats the ??'s

OpenStudy (anonymous):

@theoryofmars

OpenStudy (anonymous):

sorry it wasnt loading for me

OpenStudy (anonymous):

The diagonal of a square has a length of \[8\sqrt{2}\]cm. Find the length of a side of the square. a.)8cm b.)4cm c.)\[2\sqrt{2}cm\] d.)\[5\sqrt{2}cm\]

OpenStudy (anonymous):

@nateisboss

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Write the expression in simplified radical form. \[\sqrt{50}\] a.)\[2\sqrt{5}\] b.)\[10\sqrt{5}\] c.)\[5\sqrt{2}\] d.)5

OpenStudy (anonymous):

b

OpenStudy (anonymous):

can you put like the answer beside the letter bc my comment section is just wacky atm

OpenStudy (anonymous):

1.b.)4cm 2.b.)10√5

OpenStudy (anonymous):

b

OpenStudy (anonymous):

no its b.) 10√5

OpenStudy (anonymous):

thats what i said

OpenStudy (anonymous):

Simplify completely: